# Improving Genomic Prediction using High-dimensional Secondary Phenotypes: the Genetic Latent Factor Approach

### v1

@@ Abstract.
Decreasing costs and new technologies have led to an increase in the amount of data available to plant breeding programs. High-throughput phenotyping (HTP) platforms routinely generate high-dimensional datasets of secondary features that may be used to improve genomic prediction accuracy. However, integration of this data comes with challenges such as multicollinearity, parameter estimation in $p>n$ settings, and the computational complexity of many standard approaches. Several methods have emerged to analyze such data, but interpretation of model parameters often remains challenging.
We propose genetic factor best linear unbiased prediction (gfBLUP), a seven-step prediction pipeline that reduces the dimensionality of the original secondary HTP data using generative factor analysis. In short, gfBLUP uses redundancy filtered and regularized genetic and residual correlation matrices to fit a maximum likelihood factor model and estimate genetic latent factor scores. These latent factors are subsequently used in multi-trait genomic prediction. Our approach performs on par or better than alternatives in extensive simulations and a real-world application, while producing easily interpretable and biologically relevant parameters. We discuss several possible extensions and highlight gfBLUP as the basis for a flexible and modular multi-trait genomic prediction framework.
@@ 1. Introduction
Genomic prediction has revolutionized plant breeding programs over the past two decades. Its use allows maintaining and evaluating breeding populations larger than ever before. As a result, breeding programs have seen increases in efficiency leading to more rapid development of varieties that meet todays challenges [1]. Genomic prediction is a powerful tool in ensuring food security in the face of a growing population and more extreme climate conditions [2].
The practice of genomic prediction involves training a statistical model using evaluated genotypes, followed by making trait predictions for untrialed test genotypes [3]. The model is mainly used in a univariate form (gBLUP, genomic best linear unbiased prediction) where the predictions are sums of genomic marker effects [4]. These predictions allow breeders to make selections without the need for extensive field trials.
Recent years have seen an increase in the use of high-throughput phenotyping (HTP) technologies resulting in the characterization of individuals in more aspects than just those of the genomic and trait phenotypic data. Examples include the collection of hyperspectral reflectivity measurements on the crop canopy in large field trials, metabolic profiling of plant leaves, and gene expression profiling using DNA microarrays [e.g., 5]. These secondary phenotypes can be collected for either the training set alone (termed CV1, cross-validation scenario 1), or both the training and test individuals (CV2, cross-validation scenario 2) [6]. If of sufficient quality, the secondary data can contain valuable information that readily improves genomic predictions for the trait of interest (focal trait). There are several such examples where integration of genomic and phenomic data leads to higher accuracies [7]
Univariate genomic prediction uses only genomic information and focal trait information to obtain predictions. To integrate phenomics and genomics, the univariate model can be extended to a multivariate form relatively easily by modelling the focal trait as well as the secondary phenotypes as functions of the SNP markers [8]. Direct integration of HTP data to improve predictions is challenging for several reasons, however. First of all, secondary HTP data is often high-dimensional complicating parameter estimation. Second, subsets of secondary features are often highly correlated. This multicollinearity can greatly complicate the interpretation of model parameters [9]. Finally, even if the number of observations were greater than the number of features, and features were not highly correlated, direct integration would still be computationally demanding as the calculation of the BLUPs requires inverting a symmetric matrix whose dimension equals the number of genotypes times the number of features.
@@ 1.1. Current methodology
A number of methods have been developed to deal with the high dimensionality, multicollinearity, and computational complexity. Lopez-Cruz et al. [11] proposed linearly combining the secondary features into a single regularized selection index (siBLUP), thereby reducing the dimensionality to one. The lsBLUP method introduced by Arouisse et al. [8] is similar in the sense that it reduces the dimensionality to one, but uses a LASSO prediction of the focal trait using the HTP features instead of a regularized selection index. Another alternative is MegaLMM as proposed by Runcie et al. [10]. MegaLMM constructs an arbitrary number of independent factors and errors that are linear combinations of the original features. Finally, neural networks and deep learning have also been investigated as options for the integration of secondary HTP and genomic data [12]. Due to their flexible nature they present an attractive option to model complex interactions between many features.
@@ 1.2. The proposed model & overview
We introduce gfBLUP (genetic factor BLUP) to deal with the challenges of secondary HTP data. It is based on the idea that HTP features typically form a set of many noisy measurements of a much lower dimensional set of latent features [13]. In short, we use factor analysis to reduce the dimensionality of the HTP data to a data-driven number of uncorrelated latent factors which are subsequently used in multivariate genomic prediction. We avoid any manual hyperparameter tuning by using a data-driven approach to determine the dimensionality of the factor model. Our approach is also flexible in the sense that it can accommodate latent structures of any dimension lower than the well-known Ledermann bound [14]. This flexibility means that complex structures underlying high-dimensional data can be accurately modeled. Furthermore, the factor loadings and genetic correlations between factors and the focal trait are interpretable and reveal the relevant biological patterns in the otherwise noisy HTP data.
The remainder of this paper is organized as follows. We elaborate the proposed factor analytic and multivariate genomic prediction approach in Section 2, showing how a seven-step prediction pipeline produces predicted breeding values as well as interpretable parameters. Section 3 provides details and results on a comprehensive simulation study comparing the performances of different methods as well as the importance of various genetic parameters. In Section 4 we apply gfBLUP and alternatives to a wheat yield and hyperspectral reflectivities dataset to evaluate their real-world performance and show how gfBLUP produces biologically relevant and interpretable parameters. Finally, we discuss several possible extensions to the proposed approach in Section 5. Information on notation can be found in Section S.1 of the Supplementary Material (SM).
@@ 2. Model
2.1. Model and assumptions
Consider the full phenotypic data matrix $\mathbf{Y}_{sf}\in\mathbb{R}^{n\times(p+1)}$ of the secondary features and focal trait. We assume a balanced design for simplicity, but note that this is not required. In a balanced design $n=n_{g}r$, i.e., the number of observations is equal to the number of genotypes multiplied by the number of replicates per genotype. The matrix $\mathbf{Y}_{sf}$ can be partitioned:
(1) $\mathbf{Y}_{sf}=\begin{bmatrix}\mathbf{Y}_{s}&\mathbf{y}_{f}\end{bmatrix}\text{,}$
where the matrix $\mathbf{Y}_{s}$ is the $n\times p$ matrix containing the mean-centered secondary features and $\mathbf{y}_{f}$ is the $n\times 1$ column vector containing the mean-centered focal trait. Assuming no covariates, all phenotypes are composed of genetic and residual components:
$\mathbf{Y}_{sf}=\mathbf{G}_{sf}+\mathbf{E}_{sf}\text{.}$
@@ 2.1. Model and assumptions
$\operatorname{vec}({\mathbf{Y}_{sf}})=\operatorname{vec}({\mathbf{G}_{sf}})+\operatorname{vec}({\mathbf{E}_{sf}})\sim\mathcal{N}_{n(p+1)}(\mathbf{0},\bm{\Sigma}^{g}\otimes\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}+\bm{\Sigma}^{\epsilon}\otimes\mathbf{I}_{n})\text{,}$
where $\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}$ and $\mathbf{I}_{n}$ are the row covariances of $\mathbf{G}_{sf}$ and $\mathbf{E}_{sf}$, respectively. The $n\times n_{g}$ incidence matrix $\mathbf{Z}$ links individuals to genotypes. Similarly, $\bm{\Sigma}^{g}$ and $\bm{\Sigma}^{\epsilon}$ are the column covariances of $\mathbf{G}_{sf}$ and $\mathbf{E}_{sf}$. These covariance matrices can be block-partitioned:
$\begin{split}\bm{\Sigma}^{g}&=\begin{pmatrix}\bm{\Sigma}_{ss}^{g}&\bm{\sigma}_{sf}^{g}\\ \bm{\sigma}_{fs}^{g}&\sigma_{ff}^{g}\\ \end{pmatrix}=\begin{pmatrix}\bm{\Sigma}_{s.}^{g}\\ \bm{\sigma}_{f.}^{g}\end{pmatrix}\\ \bm{\Sigma}^{\epsilon}&=\begin{pmatrix}\bm{\Sigma}_{ss}^{\epsilon}&\bm{\sigma}_{sf}^{\epsilon}\\ \bm{\sigma}_{fs}^{\epsilon}&\sigma_{ff}^{\epsilon}\\ \end{pmatrix}=\begin{pmatrix}\bm{\Sigma}_{s.}^{\epsilon}\\ \bm{\sigma}_{f.}^{\epsilon}\end{pmatrix}\text{,}\end{split}$
where $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$ contain the genetic and residual covariances among secondary features, $\sigma_{ff}^{g}$ and $\sigma_{ff}^{\epsilon}$ are the scalar focal trait genetic and residual variances, and $\bm{\sigma}_{sf}^{g}=(\bm{\sigma}_{fs}^{g})^{\top}$ and $\bm{\sigma}_{sf}^{\epsilon}=(\bm{\sigma}_{fs}^{\epsilon})^{\top}$ contain the genetic and residual covariances between secondary features and the focal trait. The method we propose is initially unsupervised as we first focus on the secondary feature data matrix $\mathbf{Y}_{s}$ and the corresponding covariance matrices $\bm{\Sigma}_{ss}^{g}$ and $\bm{\Sigma}_{ss}^{\epsilon}$. We then use focal trait training data to train the prediction model.
While we require plot-level data to estimate covariance matrices, genomic prediction typically uses best linear unbiased estimates (BLUEs). For a completely randomized design we can obtain these BLUEs for a single genotype $j$ by simply taking the genotypic mean of all replicates $i=1,\dots,r$ [15]:
@@
Note that as the residual vectors $\bm{\epsilon}_{s(i|j)}$ of the replicates of genotype $j$ are iid random vectors, basic covariance algebra shows that the covariance of $\bar{\bm{\epsilon}}_{s(j)}$ is simply $r^{-1}\bm{\Sigma}_{ss}^{\epsilon}$ [15] (see also SM Section S.3). As a result, $\bar{\mathbf{y}}_{s(j)}\sim\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma}_{ss}^{p})$, where $\bm{\Sigma}_{ss}^{p}=\bm{\Sigma}_{ss}^{g}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon}$, assuming for a moment that $\mathbf{K}\circ\mathbf{I}_{n_{g}}=\mathbf{I}_{n_{g}}$.
In gfBLUP, we use a factor analytic structure for $\bm{\Sigma}_{ss}^{g}$. This follows from the assumption that the genetic components $\mathbf{g}_{s(j)}$ can be written as a weighted sum of $m$ latent factors and $p$ errors:
@@
where $\mathbf{\Lambda}$ is a $p\times m$ matrix containing factor loadings, $\bm{\xi}_{j}$ is an $m\times 1$ vector containing the latent factor scores for genotype $j$, and $\bm{\varepsilon}_{j}$ contains the factor model errors. Substituting (3) into (2) gives us the full model for the phenotypic data:

### v3

@@ Abstract.
Decreasing costs and new technologies have led to an increase in the amount of data available to plant breeding programs. High-throughput phenotyping (HTP) platforms routinely generate high-dimensional datasets of secondary features that may be used to improve genomic prediction accuracy. However, integration of these data comes with challenges such as multicollinearity, parameter estimation in $p>n$ settings, and the computational complexity of many standard approaches. Several methods have emerged to analyze such data, but interpretation of model parameters often remains challenging. We propose genetic latent factor best linear unbiased prediction (glfBLUP), a prediction pipeline that reduces the dimensionality of the original secondary HTP data using generative factor analysis. In short, glfBLUP uses redundancy filtered and regularized genetic and residual correlation matrices to fit a maximum likelihood factor model and estimate genetic latent factor scores. These latent factors are subsequently used in multi-trait genomic prediction. Our approach performs better than alternatives in extensive simulations and a real-world application, while producing easily interpretable and biologically relevant parameters. We discuss several possible extensions and highlight glfBLUP as the basis for a flexible and modular multi-trait genomic prediction framework.
@@ 1. Introduction
Genomic prediction has revolutionized plant breeding programs over the past two decades. Its use allows maintaining and evaluating breeding populations larger than ever before. As a result, breeding programs have seen increases in efficiency leading to more rapid development of varieties that meet today’s challenges (Crossa et al., 2013). Genomic prediction is a powerful tool in ensuring food security in the face of a growing population and more extreme climate conditions (Harfouche et al., 2019).
The practice of genomic prediction involves training a statistical model using evaluated genotypes, followed by making trait predictions for untrialed test genotypes (Meuwissen et al., 2001). The model is mainly used in a univariate form (gBLUP, genomic best linear unbiased prediction) where the predictions are sums of genomic marker effects (Endelman, 2011). These predictions allow breeders to make selections without the need for extensive field trials.
Recent years have seen an increase in the use of high-throughput phenotyping (HTP) technologies resulting in the characterization of individuals in more aspects than just those of the genomic and trait phenotypic data. Examples include the collection of hyperspectral reflectivity measurements on the crop canopy in large field trials, metabolic profiling of plant leaves, and gene expression profiling using DNA microarrays (e.g., Krause et al., 2019). These secondary phenotypes can be collected for either the training set alone (termed CV1, cross-validation scenario 1), or both the training and test individuals (CV2, cross-validation scenario 2) (Runcie and Cheng, 2019). If of sufficient quality, the secondary data can contain valuable information that readily improves genomic predictions for the trait of interest (focal trait). There are several such examples where integration of genomic and phenomic data leads to higher accuracies (Adak et al., 2023)
Univariate genomic prediction uses only genomic information and focal trait information to obtain predictions. To integrate phenomics and genomics, the univariate model can be extended to a multivariate form relatively easily by modeling the focal trait as well as the secondary phenotypes as functions of the SNP markers (Arouisse et al., 2021). Direct integration of HTP data to improve predictions is challenging for several reasons, however. First of all, secondary HTP data are often high-dimensional complicating parameter estimation. Second, subsets of secondary features are often highly correlated. This multicollinearity can greatly complicate the interpretation of model parameters (Fotheringham and Oshan, 2016). Finally, even if the number of observations were greater than the number of features, and features were not highly correlated, direct integration would still be computationally demanding as the calculation of the BLUPs requires inverting a symmetric matrix whose dimension equals the number of genotypes times the number of features.
@@ 1.1. Current methodology
A number of methods have been developed to deal with the high dimensionality, multicollinearity, and computational complexity. Lopez-Cruz et al. (2020) proposed linearly combining the secondary features into a single regularized selection index (siBLUP), thereby reducing the dimensionality to one. The lsBLUP method introduced by Arouisse et al. (2021) is similar in the sense that it reduces the dimensionality to one, but uses a LASSO prediction of the focal trait using the HTP features instead of a regularized selection index. Another alternative is MegaLMM as proposed by Runcie et al. (2021). MegaLMM constructs an arbitrary number of independent factors and unique trait contributions that are linear combinations of the original features. Neural networks and deep learning have also been investigated as options for the integration of secondary HTP and genomic data (Togninalli et al., 2023). Due to their flexible nature they present an attractive option to model complex interactions between many features.
@@ 1.2. The proposed model & overview
We introduce glfBLUP (genetic latent factor BLUP) to deal with the challenges of secondary HTP data. It is based on the idea that HTP features typically form a set of many noisy measurements of a much lower dimensional set of latent features (Münch et al., 2022). In short, we use factor analysis to reduce the dimensionality of the HTP data to a data-driven number of uncorrelated latent factors which are subsequently used in multivariate genomic prediction. We avoid any manual hyperparameter tuning by using a data-driven approach to determine the dimensionality of the factor model. Our approach is also flexible in the sense that it can accommodate latent structures of any dimension lower than the Ledermann bound (Ledermann, 1937). This flexibility means that complex structures underlying high-dimensional data can be accurately modeled. Furthermore, the factor loadings and genetic correlations between factors and the focal trait are interpretable and reveal the relevant biological patterns in the otherwise noisy HTP data.
The remainder of this paper is organized as follows. We elaborate the used notation and the proposed factor analytic multivariate genomic prediction approach in Section 2. Section 3 provides details and results on a comprehensive simulation study comparing the performances of different methods as well as the importance of various genetic parameters. In Section 4 we apply glfBLUP and alternatives to a wheat yield and hyperspectral reflectivities dataset to evaluate their real-world performance and show how glfBLUP produces biologically relevant and interpretable parameters. Finally, we discuss several possible extensions to the proposed approach in Section 5.
@@ 2. Model
2.1. Notation
We use $\circ$ and $\otimes$ to denote the Hadamard and Kronecker products. We use $diag(s_{11},\dots,s_{pp})$ to refer to a $p$-dimensional diagonal matrix with elements $s_{11},\dots,s_{pp}$. We denote the column-wise vectorization of a matrix by $\operatorname{vec}(\mathbf{A})$ and use $\operatorname{vec}^{-1}(\mathbf{A})$ for the reverse operation. The cardinality of a set $\mathcal{A}$ is denoted by $\text{card}(\mathcal{A})$. Some further information and examples on the partitioning of data and covariance matrices, as well as some general background information on genomic prediction, can be found in SM Section 1.
2.2. Modeling of secondary features
The pipeline we propose uses a fully unsupervised dimensionality reduction before including factor scores as additional traits in multi-variate genomic prediction. This section contains details on the steps involved in the dimensionality reduction. Section 2.3 then provides some details on the multivariate genomic prediction.
2.2.1. Model and assumptions
Consider the secondary feature phenotypic data matrix $\mathbf{Y}_{s}\in\mathbb{R}^{n\times p}$. We assume a balanced design for simplicity, but note that this is not required. In a balanced design $n=n_{g}r$, i.e., the number of observations is equal to the number of genotypes multiplied by the number of replicates per genotype. Assuming no covariates, all secondary phenotypes are composed of genetic and residual components:
$\mathbf{Y}_{s}=\mathbf{G}_{s}+\mathbf{E}_{s}\text{.}$
@@ 2.2.1. Model and assumptions
$\operatorname{vec}({\mathbf{Y}_{s}})=\operatorname{vec}({\mathbf{G}_{s}})+\operatorname{vec}({\mathbf{E}_{s}})\sim\mathcal{N}_{np}(\mathbf{0},\bm{\Sigma}^{g}_{ss}\otimes\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}+\bm{\Sigma}^{\epsilon}_{ss}\otimes\mathbf{I}_{n})\text{,}$
where $\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}$ and $\mathbf{I}_{n}$ are the row covariances of $\mathbf{G}_{s}$ and $\mathbf{E}_{s}$, respectively. The $n\times n_{g}$ incidence matrix $\mathbf{Z}$ links individuals to genotypes. Similarly, $\bm{\Sigma}^{g}_{ss}$ and $\bm{\Sigma}^{\epsilon}_{ss}$ are the column covariances of $\mathbf{G}_{s}$ and $\mathbf{E}_{s}$. We use the subscript ${ss}$ to denote that $\bm{\Sigma}^{g}_{ss}$ and $\bm{\Sigma}^{\epsilon}_{ss}$ contain genetic and residual covariances among secondary features.
While we require plot-level data to estimate covariance matrices, two-stage genomic prediction typically uses best linear unbiased estimates (BLUEs). For a completely randomized design we can obtain these BLUEs for a single genotype $j$ by simply taking the genotypic mean of all replicates $i=1,\dots,r$ (Kruijer et al., 2020):
@@
Note that as the residual vectors $\bm{\epsilon}_{s(i|j)}$ of the replicates of genotype $j$ are iid random vectors, basic covariance algebra shows that the covariance of $\bar{\bm{\epsilon}}_{s(j)}$ is simply $r^{-1}\bm{\Sigma}_{ss}^{\epsilon}$ (Kruijer et al., 2020) (see also Supplementary Material (SM) Section 2). As a result, $\bar{\mathbf{y}}_{s(j)}\sim\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma}_{ss}^{p})$, where $\bm{\Sigma}_{ss}^{p}=\bm{\Sigma}_{ss}^{g}+r^{-1}\bm{\Sigma}_{ss}^{\epsilon}$, assuming for a moment that $\mathbf{K}\circ\mathbf{I}_{n_{g}}=\mathbf{I}_{n_{g}}$. Also note that the assumption of iid errors is valid only if data is truly from a completely randomized design. If some pre-processing is applied in a two-stage modeling approach, e.g., spatial correction, the assumption is an approximation only.
In glfBLUP, we use a factor analytic structure for $\bm{\Sigma}_{ss}^{g}$. This follows from the assumption that the genetic components $\mathbf{g}_{s(j)}$ can be written as a weighted sum of $m$ latent factors and $p$ factor model errors:
@@
where $\mathbf{\Lambda}$ is a $p\times m$ matrix containing factor loadings, $\bm{\xi}_{j}$ is an $m\times 1$ vector containing the latent factor scores for genotype $j$, and $\bm{\varepsilon}_{j}$ contains the factor model errors. Substituting (2) into (1) gives us the full model for the secondary phenotypic data: